OpenStudy (anonymous):

What is the slope of the line? y= -3x+8

jimthompson5910 (jim_thompson5910):

\[\Large y = -3x+8\] is in the form \[\Large y = mx+b\]

jimthompson5910 (jim_thompson5910):

notice how the -3 and m match up \[\Large y = \color{red}{-3}x+8\] \[\Large y = \color{red}{m}x+b\]

OpenStudy (ram231):

so 8 is known as the y intercept right?

jimthompson5910 (jim_thompson5910):

m = slope b = y intercept

OpenStudy (seeing.in.pixels):

answer: -3